Insurance companies are interested in knowing the population percent of drivers who always buckle up before riding in a car. They randomly survey 393 drivers and find that 305 claim to always buckle up. Construct a 87% confidence interval for the population proportion that claim to always buckle up. Use interval notation, for example, [1,5]
Answer #1

n = 393 , X = 305

87% confidence level

z = 0.8251

p^= X/n = 305/393 = 0.77608

(p^- z * sqrt(pq/n) , p^+ z* sqrt(pq/n))

=(0.77608 - 0.8251*sqrt(0.77608 * (1-0.77608 ) / 393) , 0.77608 + 0.8251 * sqrt(0.77608 * (1-0.77608 ) / 393) )

= ( 0.75873 , 0.79343)

Please rate